# Large temperature-up-jump simulations of a binary Lennard-Jones system

### v1

@@ IV Validating the triangular relation in simulations and determining the material time
    The incoherent intermediate scattering function $F_{s}(t_{1},t_{2})$ calculated from the particle displacement from time $t_{1}$ to time $t_{2}$, evaluated at the wavevector corresponding to the first maximum of the radial distribution function (we determined this at the final temperature $0.48$ but the peak position was virtually constant throughout all simulations),
@@
    $\chi_{4}(t_{1},t_{2},q)=N_{A}\Biggl[\Bigg\langle\left(\frac{1}{N_{A}}\sum_{i=1}^{N_{A}}\cos({\bf q}\cdot({\bf r}_{i}(t_{2})-{\bf r}_{i}(t_{1})\right)^{2}\Bigg\rangle-F_{s}^{2}(t_{1},t_{2})\Biggr]\,.$ (7)
    The length of $\bf q$} is the same as above for the incoherent intermediate scattering function;
@@
    The non-Gaussian parameter $\alpha_{2}(t_{1},t_{2})$ (which is zero for Gaussian diffusion) defined by
@@
Figure 4: Left Two-time functions plotted during aging as functions of the time difference $t_{2}-t_{1}$. The left column gives data for the small up jump, the right for the large one. The color of the curves vary from blue for small values of $t_{1}$ to red as this value increases. The following two-time observables are considered as functions of $t_{2}-t_{1}$: (a-b) Per-particle potential energy autocorrelation function; (c-d) Self part of the intermediate scattering function; (e-f) Mean-square displacement; (g-h) Dynamical susceptibility (Eq. (7)); (i-j) Non-Gaussian parameter (Eq. (8)). The curves are identical at large $t_{1}$ for both jumps because they share the same final temperature while they differ at small $t_{1}$s. Right Same data plotted as functions of the material-time difference $\xi_{2}-\xi_{1}$. There is a significantly better collapse for the small up jump than for the large one.
@@
Figure 4 reports the results of our simulations for the temperature up jumps from $T=0.43$ and $T=0.37$ to $T=0.48$. The left two columns show the time-autocorrelation functions during aging plotted as a function of $t_{2}-t_{1}$ with the smaller jump to the left, parametrized by the waiting time $t_{1}$. As expected, curve collapse is not observed for any of the five quantities. Instead a general speed up is seen, with the short-waiting time bluish curves moving to shorter times. This is the well-known “self-catalyzed” behavior, reflecting the fact that the structure gradually approaches that of equilibrium at $T=0.48$ at which relaxation is much faster than at the two starting temperatures. Comparing the results for the jump from $T=0.43$ with those of the $T=0.37$ to $T=0.48$ jump, there is a considerably larger spread in the latter case. This reflects that fact that the relaxation rates vary more, the larger the difference is between two temperatures. At the longest times corresponding to the red curves, a collapse is observed because equilibrium at $T=0.48$ is approached.
@@
The right two columns of Fig. 4 give the main results of this paper, investigating to which extent the material time collapses all the autocorrelation functions to be unique functions of $\xi_{2}-\xi_{1}$ in which $\xi_{j}$ is the value of the material time at time $t_{j}$. There is no collapse at short material times because, as already mentioned, the material-time concept makes little sense on the short time scale corresponding to phonon times. At larger values of $\xi_{2}-\xi_{1}$ there is a much better collapse, but to a varying degree, and curiously, bluish curves corresponding to short waiting time tend to collapse better then intermediate values of $t_{1}$. Generally, the small temperature jump results in a better collapse than the large one.
@@
Interestingly, the different quantities do not show the same behavior. Starting with the small jump, the potential-energy autocorrelation function and the incoherent intermediate scattering function show quite good collapse, while the mean-square displacement, $\chi_{4}$, and $\alpha_{2}$, do not collapse (though much more so than in the real-time panel). The situation is worse for the large up jump in which case none of the five quantities exhibit a convincing collapse. Even here, however, most of the huge variation of the real-time panel is eliminated by the introduction of the material time. One also observes that while the $\chi_{4}$ curves systematically shift toward later times compared to the final equilibrium value, the $\alpha_{2}$ curves behave oppositely: those at small $t_{1}$ relax faster than the curves at larger $t_{1}$.
@@ VI Discussion
We have seen that the triangular relation is well obeyed for the potential energy. Despite this, the material time defined from the potential energy is not able to describe well the aging of the five autocorrelation functions following, in particular, the largest temperature up jump. Two possible explanations of this comes to mind. One possibility is that each quantity has its own material time in the sense that its aging is much better described in this way than by using a common material time. Another possibility is that dynamic heterogeneities invalidates the existence of any global material time in large up jumps, as suggested some time ago by Castillo and Parsaeian [31]. This would make sense physically because little relaxation can take place in a slow region. Indeed, for a two-dimensional system Herrero et al. have recently demonstrated the existence of growing regions of activity during large temperature up jumps [32]. On the other hand, even equilibrium highly viscous liquids exhibit dynamic heterogeneities, and here there is time-translational invariance of all time-autocorrelation functions and no need for introducing local times. Thus if local clocks are needed to describe large up jumps, this can only be the case under these specific aging conditions. It appears that more work is needed to clarify whether the introduction of local clocks somehow can “save” the TN formalism or it simply has to be discarded for large temperature up jumps.

### v2

@@ IV Validating the triangular relation and determining the material time
    The incoherent intermediate scattering function $F_{s}(t_{1},t_{2})$ calculated from the particle displacement from time $t_{1}$ to time $t_{2}$, evaluated at the wavevector corresponding to the first maximum of the radial distribution function determined at the final temperature $0.48$ (the peak position was virtually constant throughout all simulations),
@@
    $\chi_{4}(t_{1},t_{2},q)=N_{A}\Biggl[\Bigg\langle\left(\frac{1}{N_{A}}\sum_{i=1}^{N_{A}}\cos({\bf q}\cdot({\bf r}_{i}(t_{2})-{\bf r}_{i}(t_{1})\right)^{2}\Bigg\rangle-F_{s}^{2}(t_{1},t_{2})\Biggr]\,;$ (7)
@@
    The non-Gaussian parameter $\alpha_{2}(t_{1},t_{2})$ (which is zero for Gaussian diffusion),
@@
Figure 4: Left Two-time functions plotted during aging as functions of the time difference $t_{2}-t_{1}$. The left column gives data for the smaller up jump, the right for the larger one. The color of the curves vary from blue for small values of $t_{1}$ to red as this value increases. The following two-time observables are considered as functions of $t_{2}-t_{1}$: (a-b) Per-particle potential energy time-autocorrelation function; (c-d) Self part of the intermediate scattering function; (e-f) Mean-square displacement; (g-h) Dynamical susceptibility; (i-j) Non-Gaussian parameter. The curves are identical at large $t_{1}$ for both jumps because they share the same final temperature while they differ at small $t_{1}$. Right Same data plotted as functions of the material-time difference $\xi_{2}-\xi_{1}$, this time with peak normalisation of $\chi_{4}$ and $\alpha_{2}$. There is a significantly better collapse for the smaller up jump than for the larger one.
@@
Figure 4 reports the results of our simulations for temperature up jumps from $T=0.43$ and $T=0.37$ to $T=0.48$. The left two columns show the time-autocorrelation functions during aging plotted as functions of $t_{2}-t_{1}$ with the smaller jump to the left, parametrized by the waiting time $t_{1}$. As expected, curve collapse is not observed for any of the five quantities. Instead a general speed up is seen, with the short-waiting time (bluish) curves moving to shorter times. This is the well-known “self-catalyzed” behavior, reflecting the fact that the structure gradually approaches that of equilibrium at $T=0.48$ at which relaxation is much faster than at the two starting temperatures. Comparing the results for the jump from $T=0.43$ to those of the $T=0.37$ to $T=0.48$ jump, there is a considerably larger spread in the latter case. This reflects that fact that the relaxation rates vary more, the larger the difference is between the initial and final temperatures of a jump. At the longest times corresponding to the red curves, a beginning collapse is observed because equilibrium at $T=0.48$ is approached.
@@
The right two columns of Fig. 4 give the main results of this paper, investigating to which extent the material time defined from the potential-energy time-autocorrelation collapses all the two times functions to be unique functions of $\xi_{2}-\xi_{1}$ in which $\xi_{j}$ is the value of the material time at time $t_{j}$. Note that here the non-monotonic quantities, the dynamical susceptibility $\chi_{4}$ and the non-Gaussian parameter $\alpha_{2}$, are normalised by their peak values; this is because we are only interested in the evolution of the shape of two-time correlation during aging. There is no collapse at short material times because, as already mentioned, the material-time concept makes little sense on the short time scale corresponding to phonon times and slightly longer times. At larger values of $\xi_{2}-\xi_{1}$ there is a much better collapse, though to a varying degree and, curiously, the bluish curves corresponding to short waiting time collapse better than the curves corresponding to intermediate values of $t_{1}$. Generally, the smaller temperature jump results in a better collapse than the larger one.
@@
Interestingly, the different quantities do not show the same behavior. Starting with the smaller jump, all quantities show quite good collapse, apart from $\chi_{4}$ and $\alpha_{2}$ that do not collapse well (though much better so than in the real-time panel). The situation is worse for the large up jump in which case none of the five quantities exhibit a really convincing collapse. Even here, however, most of the huge variation of the real-time panel is eliminated by the introduction of the material time. For the large jump, even the potential‑energy autocorrelation displays small slope differences preventing full collapse, indicating that a global clock might no longer be defined. One also observes that while the $\chi_{4}$ curves systematically shift toward later times compared to the final equilibrium value, the $\alpha_{2}$ curves behave oppositely: those at small $t_{1}$ relax faster than those at larger $t_{1}$. $\alpha_{2}$ probes the appearance of flow events that correspond to large deviations from the Gaussian dynamics, while the self part of the dynamical susceptibility reflects when those events become spacially correlated [34, 33]. One may thus speculate that the difference in behavior between $\chi_{4}$ and $\alpha_{2}$ might reflect heterogeneity occurring before cooperativity in the particles dynamics.
@@ VI Discussion
We have shown that the triangular relation is well obeyed for the potential energy. Despite this, the material time defined from the potential energy is not able to describe well the aging of the five autocorrelation functions following very large temperature up jumps. Two possible explanations of this observation comes to mind. One possibility is that each quantity has its own material time in the sense that aging is better described in this way than by using a common material time. Another possibility is that dynamic heterogeneities invalidate the existence of any global material time for large up jumps, as suggested some time ago by Castillo and Parsaeian [35]. This would make sense physically because little relaxation can take place in a slow region. Indeed, for vapour deposited ultra-stable glasses Vila-Costa et al. have recently observed the existence of growing regions of activity during large temperature up jumps [17, 36], and this observation has been reproduced in simulations [37]. On the other hand, even equilibrium highly viscous liquids exhibit dynamic heterogeneities, and here there is time-translational invariance of all time-autocorrelation functions and no need to introduce local times. Thus if local clocks are needed to describe large temperature up jumps, this can only be the case under specific aging conditions. More work is needed to clarify whether the introduction of local clocks can save the TN formalism or it must be discarded altogether for large temperature up jumps.